OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x + y = 15 4x + 3y = 38

OpenStudy (deez):

x = 15 - y 4 (15-y) + 3y = 38 60 - 4y + 3y = 38 60 - 38 = y = 22 --> x = 15 - y = -7 (x,y) = (-7 , 22)